OpenStudy (anonymous):

Suppose you have a part time job your boss offers to pay you $2 the first day with the amount to double each day how much will you be paid for the tenth day make a plan find the amounts you will be paid on the second and thrid, and fourth write each amount using exponents carry out the pla on each day you will earn __ dollars which can be written using exponents as 2 __ identify the pattern

OpenStudy (anonymous):

2, 4, 8, 16, 32, 64, 128, 256, 51, 1024. Grand total of 2046

OpenStudy (anonymous):

use sum of first-n terms of geometric sequences formula S=a(1-r^n)/(1-r) r=2 a=2 2(1-2^10)/(1-2)=2046